# WhyFlow: Interrogative Debugger for Sensemaking Taint Analysis

### v2

@@ Abstract.
Taint analysis is a security analysis technique used to track the flow of potentially dangerous data through an application and its dependent libraries. Investigating why certain unexpected flows appear and why expected flows are missing is an important sensemaking process during end-user taint analysis. Existing taint analysis tools often do not provide this end-user debugging capability, where developers can ask why, why-not, and what-if questions about dataflows and reason about the impact of configuring sources and sinks, and models of 3rd-party libraries that abstract permissible and impermissible data flows. Furthermore, a tree-view or a list-view used in existing taint-analyzer’s visualization makes it difficult to reason about the global impact on connectivity between multiple sources and sinks.
Inspired by the insight that sensemaking tool-generated results can be significantly improved by a QA inquiry process, we propose WhyFlow, a first end-user question-answer style debugging interface for taint analysis. It enables a user to ask why, why-not, and what-if questions to investigate the existence of suspicious flows, the non-existence of expected flows, and the global impact of third-party library models. WhyFlow performs speculative what-if analysis, to help a user in debugging how different connectivity assumptions affect overall results. A user study with 12 participants shows that participants using WhyFlow achieved 21% higher accuracy on average, compared to CodeQL. They also reported a 45% reduction in mental demand (NASA-TLX) and rated higher confidence in identifying relevant flows using WhyFlow. This shows WhyFlow’s potential to significantly reduce sensemaking effort.
copyright: acmlicensedjournalyear: 2026conference: IEEE/ACM International Conference on Software Engineering; April 15–17, 2026; Rio de Janeiro, Brazil
@@ 1. Introduction
To prevent security vulnerabilities, developers use taint analysis. This technique tracks potentially dangerous data as it moves through a program. Data from untrusted sources, like user input, is marked as "tainted" The analysis then monitors how this tainted data spreads. If it reaches a critical point in the code, known as a "taint sink," a warning is generated. Tainted data can be made safe by "sanitizers," which are functions that clean or encode the data. For example, the java.net.URLEncoder.encode(String input, String encoding) function escapes potentially harmful characters, preventing security issues. Since analyzing the entire program is usually impossible, practical taint analysis requires configuring simplified models of external libraries (Chibotaru2019, ). These models abstract the behavior of third-party libraries, allowing an analysis to track taint flow without analyzing the libraries’ internals.
@@
Taint analysis tools often come with default configurations, including pre-built models of external libraries. However, these default models, often automatically generated, can be inaccurate. This is because they rely on assumptions, which may be incorrect, about how libraries handle data flow. Incorrect configurations can lead to significant problems. For example, if a model incorrectly allows tainted data to flow through a library function that should sanitize it, the analysis will produce unexpected taint flows (i.e, false positives) (Livshits2009, ; Banerjee2023, ). This means it will report potential vulnerabilities that do not actually exist, creating extra, unexpected warnings for the user. Conversely, if a model incorrectly blocks the flow of tainted data when it should not, the analysis will produce missing flows (i.e., false negatives). This means it will miss real vulnerabilities, failing to report flows expected by the user.
@@
This paper focuses on end-user debugging of taint analysis that do not match a user’s expectations, in other words, debugging an end-user configuration, as opposed to debugging a taint analysis implementation. To correct configuration errors, developers must understand the impact of configuration on tool-generated warnings. This sensemaking process involves tracing the reported dataflows, understanding why specific warnings are generated, and how these warnings relate to the models of third-party libraries.
@@
We introduce WhyFlow, a novel end-user debugger that brings interactive, question-based sensemaking to taint analysis. Drawing inspiration from interrogative debugging principles, WhyFlow enables developers to ask why, why not, and what if questions about their analysis configuration. WhyFlow provides six customizable question templates, which are then concretized by a user. A user can make sense of currently detected taint flows and hypothetical flows through interactive QA. Unlike state-of-the-art tools like CodeQL that rely on list views, WhyFlow visualizes these reported and hypothetical flows graphically, offering a more intuitive understanding of configuration changes. Below, we discuss two key features,Inquiry-based Sensemaking and Visualization of Global Impact in detail.
@@
Inquiry-based Sensemaking. To effectively debug taint analysis results, developers need to investigate why specific dataflows are reported and why others are not. Unfortunately, many tools don’t allow developers to interactively ask questions about these flows. Furthermore, to efficiently narrow down their analysis, developers must understand how modifying the models of external libraries impacts the permitted or blocked dataflows.
@@
Similarly, a user could select the why-not template, which asks, “Why is there no taint flow from a source to a sink?” They can then specify the source and sink of interest. For example, they could ask, “Which third-party library models could potentially block taint flows from the source msg:HttpRequest to the sink ErrorTypeAwareLogger.warn()?” Alternatively, a user could also select the what-if template, which asks a speculative analysis question about “which sinks would no longer be reachable if third-party library models were to be configured as a sanitizer from source to sink?”
@@
Visualization of Global Impact. Once a user initiates a query within WhyFlow, specifying their target source, sink, and any relevant external API calls, WhyFlow generates an interactive graph visualization, providing a holistic view of the taint flow. This graph, as seen in the G pane of Figure 1, visually maps the program’s taint flows as interconnected nodes, each representing a distinct stage in the taint flow. This visualization reveals the global impact of configuration choices through the network of connected nodes. To facilitate sensemaking, WhyFlow employs a color-coded scheme: source nodes are highlighted in green, sink nodes in red, and external library nodes in orange. This color differentiation draws attention to the impact of configuration choices. Furthermore, templated queries are designed to illuminate how which third-party library’s model can impact multiple sinks, multiple sources, etc, making global impact salient through visualization.
@@
User study. We conducted a within-subject study with a factorial crossover on 12 participants (graduate students and professional developers), who inspected taint-analysis warnings generated by CodeQL (avgustinov2016ql, ; szabo2023incrementalizing, ). Each participant answered eight questions per task (16 total) designed to reflect a realistic sensemaking process of tool-generated warnings. During these tasks, a user had to answer questions about the impact of third-party library models on taint analysis results, by identifying pass-through API calls, APIs that serve as sanitizers, etc. Participants increased the average question completion rate from 71% with the baseline CodeQL visualizer to 92% with WhyFlow, especially on questions that require global reasoning of multiple taint flows. When using WhyFlow, users were more accurate in identifying multiple sinks affected by the same third-party library model. For example, 50% of answers prepared with WhyFlow are correct, whereas only 17% with the baseline.
@@
We measured cognitive load via NASA-TLX questions. We found that WhyFlow led to improvement on the participants self-reported mental demand, stress, and success rate. Considering the technology acceptance model (lee2003technology, ), participants rated WhyFlow higher on both Confidence (mean 4.3 vs. 2.1) and Ease of Use (4.3 vs. 1.9) than the baseline CodeQL. Qualitative feedback highlighted that color-coded graphs and dedicated template queries helped alleviate analysis paralysis (schwartz2015paradox, ). Overall, these results underscore that WhyFlow substantially enhances users’ ability to make sense of complex taint paths and reduce their mental workload through QA-based end-user debugging.
@@
    Configuring taint analysis is tricky, often leading to unexpected results. To help users understand and fix these issues, WhyFlow is the first interactive end-user debugger that enables the investigation of how different user configuration choices in modeling third-party libraries impact the taint analysis results.
@@
    WhyFlow is equipped with template queries for ‘why’, ‘why not’, and ‘what if’ questions, which are automatically translated into logic queries; it enables users to concretize the template queries with concrete code names embedded in tool-generated warnings. WhyFlow’s graphical interface eases the examination of the global impact of models, which are difficult to reason when viewing warnings in a list view, packaged with an existing taint analyzer.
@@
    Improvement in accuracy, reduction in cognitive load, improvement in confidence and ease of use compared to CodeQL’s visualizer is statistically significant. The additional case study with two professionals corroborates this improvement in accuracy, confidence, and ease of use.
The remainder of this paper is organized as follows. Section 2 introduces a motivating example and the design goals of WhyFlow. Section 3 presents WhyFlow’s approach. Section 4 provides the study design and results from our within-subject user study. Section 5 reports the results we obtained from our user study. Section 6 discusses implications of our findings, and possible threats to validity. Section 7 presents related work. Finally, we draw the conclusions of our work in Section 8.
@@ 2. Motivation
When using modern static taint analyzers such as CodeQL (szabo2023incrementalizing, ), developers struggle to identify why unexpected flows appear or why certain expected flows are missing. These tools rely on models of external libraries, which may not always match the developer’s expectations. This paper is concerned with the problem of end-user debugging of taint analysis—i.e., understanding the impact of user configuration choices. Two key end-user debugging challenges prevent developers to easily make sense of unexpected results.
First, traditional tools lack support for effective inquiry-based debugging, presenting only isolated warnings without contextual flow connections. This forces developers into manual tracing, which hinders their ability to answer questions they have during debugging. A robust QA process, focused on sensemaking, is therefore essential to bridge this gap to enable active investigation rather than passively viewing the flow.
Second, beyond isolated flow views, developers require support for reasoning about the model’s impact on multiple dataflows. Current tools fall short, lacking the ability to easily query these complex relationships, or predict the ripple effects of model changes. This forces developers to manually piece together disparate flows, hindering them from understanding the broader consequences of modifications and leaving unreported flows completely opaque.
Prior debugging introduced WhyLine (ko2004designing, ) to let users ask “why did” or “why didn’t” on a program trace, where the context of debugging is on runtime events (e.g., runtime errors). However, “why and why-not” questions remain largely unaddressed in the context of modern taint analysis tools. Developers want customizability over taint analysis (nachtigall2022large, ), visual outputs of warnings and code (johnson2013don, ), and the ability to inspect intermediate pass through nodes of the analysis (ko2004designing, ). Yet, typical dataflow or taint analysis interfaces do not systematically address higher-level inquiries such as “Why does data from source X reach sink Y?” or “Why is no warning raised for a known bug?
@@
AffectedSinks “If we alter a third party library $Z$’s model, which sinks are affected?” “Under a new assumption of treating a third party library $Z$ as a sanitizer, which previously reported sinks are no longer reachable from XX?”
DivergentSinks “Which third party library $X$’s model could influence multiple taint flows from the same source XX?” “What are the common third-party API nodes in multiple paths originating from $X$ that split into multiple different sinks?”
DivergentSources “Which third party library $X$’s model could influence multiple taint flows reaching the same sink YY?” “What are the common third-party API nodes in multiple flow paths that eventually reach the same sink YY?”
GlobalImpact “Which third party library $Z$’s model could have the largest global influence on dataflows from $X$ to YY?” “What is the frequency of each third-party API call appearing along all paths from $X$ to $Y$ in terms of the overall usage counts?”
@@
1. WhyFlow: Why is there a taint flow from a source X to a sink Y? Suppose that Alice sees an unexpected taint-analysis warning in CodeQL from a source user.getSSN() (untrusted data) to a sink log(SSN) (potential vulnerability) (nachtigall2022large, ; johnson2013don, ; ko2004designing, ). She suspects the culprit is an imprecise third-party library model; however, manually tracing a long chain of calls is overwhelming. The WhyFlow query highlights in one view the taint flow path from a source in green to a sink in red, displaying pass-through third-party API calls in orange, shown in Figure 2. Alice then notices an intermediate API call is encrypt(SSN). She then discovers that this API is a sanitizer, and thus this warning should have not reported and encrypt()’s model should be debugged.
@@
2. WhyNotFlow: Why is there no taint flow from a source X to a sink Y? Suppose that Alice needs to investigate a missing taint flow (i.e., a bug arises at a sink, yet CodeQL does not issue a corresponding warning). She suspects that a third-party library is mistakenly modeled as a sanitizer. In Figure 3, the sensitive data travels from user.getSSN() through an external API format(SSN) to log(SSN). After investigating the intermediate flow steps, Alice finds that format(SSN) was erroneously modeled as a sanitizer.
WhyNotFlow visually identifies which APIs are acting as sanitizers (dashed arrows), pinpointing which API model could be responsible for killing a flow. WhyNotFlow performs a speculative analysis by reasoning which taint flow path could have been plausible under a configuration where a sanitizer is instead modeled as a non-sanitizer. In Figure 3, the graphical view marks the arrow after format(SSN) with a dashed line, indicating the flow currently does not exist, but would exist with a different model assumption.
@@
3. AffectedSinks: If we alter a third party library $Z$’s model, which sinks are affected? Alice wants to reason about the global impact of updating a third party library model (smith2015questions, ). Suppose that she sees an unexpected warning and considers marking a third-party library’s API as a sanitizer. However, she is concerned this update might suppress other warnings. Without WhyFlow, she would need to sift through all reported warnings, manually trace each flow from the source, and check which sinks might become unreachable after her change—a time-consuming process. Figure 4 shows how AffectedSinks automates this what-if analysis, immediately revealing all red sinks that would be “killed.”.
@@
Figure 4. AffectedSinks: “If we alter a third party library’s model, which sinks are affected?
@@
4. DivergentSinks & 5. DivergentSources: Which third party library model could influence multiple taint flows reaching the same sink (or originating from the same source)?
@@
Suppose that Alice would like to know whether a known vulnerability reaching multiple sinks could be fixed at once (smith2015questions, ; piskachev2019codebase, ). In a typical taint analyzer, it can be cumbersome to trace how a single piece of sensitive data such as an SSN propagates through multiple taint flows. Consequently, developers may not prefer to identify a common interception point instead of fixing one sink at a time. With DivergentSinks query, she can quickly locate a common point from the source that “splits” into multiple sinks. (see Figure 5).
@@
Figure 5. DivergentSinks: “Which third party library model could influence multiple taint flows from the same source?”
@@
To illustrate another scenario, imagine a single sensitive destination (sink) that receives data from multiple untrusted sources. Using the DivergentSources query, Alice can readily identify if these diverse sources converge towards the same vulnerable point. This allows her to pinpoint the specific model responsible for the convergence, i.e., the ‘culprit model’. Unlike traditional list-based warning views, WhyFlow provides a visual representation of where these multiple taint paths intersect, which makes pinpointing the culprit model significantly easier.
@@
6. GlobalImpact: Which third party library’s model could have the largest global influence on multiple flows from a source to a sink? When multiple models could explain a false or missing warning, developers often prefer a conservative fix that impacts the fewest flows (murphy2013design, ). Without WhyFlow, identifying the frequency of each API in all taint paths would require her to painstakingly review every single warning and count occurrences manually. GlobalImpact automatically computes how often each API appears across multiple paths, ranking them by frequency. In Figure 6, because extractParts(standardizedSSN) appears in more paths, it has a higher score, signaling a larger global impact.
@@
Figure 6. GlobalImpact: “Which third party library model could have the largest global influence on dataflows from a source to a sink?”

### v3

@@ Abstract.
Taint analysis is a security analysis technique used to track the flow of potentially dangerous data through an application and its dependent libraries. Investigating why certain unexpected flows appear and why expected flows are missing is an important sensemaking process during end-user taint analysis. Existing taint analysis tools often do not provide this end-user debugging capability, where developers can ask why, why-not, and what-if questions about dataflows and reason about the impact of configuring sources and sinks, and models of third-party libraries that abstract permissible and impermissible data flows. Furthermore, the tree-view or list-view used in existing taint analyzer visualizations makes it difficult to reason about the global impact on connectivity between multiple sources and sinks.
Inspired by the insight that sensemaking tool-generated results can be significantly improved by a QA inquiry process, we propose WhyFlow, the first end-user question-answer style debugging interface for taint analysis. It enables a user to ask why, why-not, and what-if questions to investigate the existence of suspicious flows, the non-existence of expected flows, and the global impact of third-party library models. WhyFlow performs speculative what-if analysis, to help a user in debugging how different connectivity assumptions affect overall results. A user study with 12 participants shows that participants using WhyFlow achieved 21% higher accuracy on average, compared to CodeQL. They also reported a 45% reduction in mental demand (NASA-TLX) and rated higher confidence in identifying relevant flows using WhyFlow. This shows WhyFlow’s potential to significantly reduce sensemaking effort.
taint analysis, end-user debugging, interrogative debugging, sensemaking, speculative analysis, human-centered static analysis
copyright: acmlicensedjournalyear: 2026conference: IEEE/ACM International Conference on Software Engineering; April 15–17, 2026; Rio de Janeiro, Brazilccs: Human-centered computing HCI design and evaluation methods
@@ 1. Introduction
To prevent security vulnerabilities, developers use taint analysis. This technique tracks potentially dangerous data as it moves through a program. Data from untrusted sources, like user input, is marked as “tainted.” The analysis then monitors how this tainted data spreads. If it reaches a critical point in the code, known as a “taint sink,” a warning is generated. Tainted data can be made safe by “sanitizers,” which are functions that clean or encode the data. For example, the java.net.URLEncoder.encode(String input, String encoding) function escapes potentially harmful characters, preventing security issues. Since analyzing the entire program is usually impossible, practical taint analysis requires configuring simplified models of external libraries (Chibotaru2019, ). These models abstract the behavior of third-party libraries, allowing an analysis to track taint flow without analyzing the libraries’ internals.
@@
Taint analysis tools often come with default configurations, including pre-built models of external libraries. However, these default models, which are often automatically generated, can be inaccurate. This is because they rely on assumptions, which may be incorrect, about how libraries handle data flow. Incorrect configurations can lead to significant problems. For example, if a model incorrectly allows tainted data to flow through a library function that should sanitize it, the analysis will produce unexpected taint flows (i.e., false positives) (Livshits2009, ; Banerjee2023, ). This means it will report potential vulnerabilities that do not actually exist, creating extra, unexpected warnings. Conversely, if a model incorrectly blocks the flow of tainted data when it should not, the analysis will produce missing flows (i.e., false negatives). This means it will miss real vulnerabilities, failing to report flows expected by the user.
@@
We introduce WhyFlow, a novel end-user debugger that brings interactive, question-based sensemaking to taint analysis. Drawing inspiration from interrogative debugging principles, WhyFlow enables developers to ask why, why not, and what if questions about their analysis configuration. This helps developers make sense of the impact of configuration and models of third-party libraries on tool-generated warnings. This sensemaking process involves tracing the reported dataflows, understanding why specific warnings are generated, and how these warnings relate to the models of third-party libraries. WhyFlow focuses on debugging an end-user configuration instead of a taint analysis implementation.
@@
WhyFlow provides six customizable question templates, which are then concretized by a user. A user can make sense of currently detected taint flows and hypothetical flows through interactive QA. Unlike state-of-the-art tools like CodeQL that rely on list views, WhyFlow visualizes these reported and hypothetical flows graphically, offering a more intuitive understanding of configuration changes. Below, we discuss two key features: Inquiry-based Sensemaking and Visualization of Global Impact in detail.
@@
Inquiry-based Sensemaking. To effectively debug taint analysis results, developers need to investigate why specific dataflows are reported and why others are not. Unfortunately, many tools do not allow developers to interactively ask questions about these flows. Furthermore, to efficiently narrow down their analysis, developers must understand how modifying the models of external libraries impacts the permitted or blocked dataflows.
@@
Similarly, a user could select the why-not template, which asks, “Why is there no taint flow from a source to a sink?” They can then specify the source and sink of interest. For example, they could ask, “Which third-party library models could potentially block taint flows from the source msg:HttpRequest to the sink ErrorTypeAwareLogger.warn()?” Alternatively, a user could also select the what-if template, which asks a speculative analysis question: “Which sinks would no longer be reachable if third-party library models were configured as a sanitizer from source to sink?”
@@
Visualization of Global Impact. Once a user initiates a query within WhyFlow, specifying their target source, sink, and any relevant external API calls, WhyFlow generates an interactive graph visualization, providing a holistic view of the taint flow. This graph, as seen in the G pane of Figure 1, visually maps the program’s taint flows as interconnected nodes, each representing a distinct stage in the taint flow. This visualization reveals the global impact of configuration choices through the network of connected nodes. To draw attention to the impact of configuration choices, WhyFlow employs a color-coded scheme: source nodes are highlighted in green, sink nodes in red, and external library nodes in orange. Templated queries are designed to illuminate how third-party library models can impact multiple sinks and multiple sources, making the global impact salient through visualization.
@@
User study: We conducted a within-subject study with a factorial crossover involving 12 participants (graduate students and professional developers) who inspected taint-analysis warnings generated by CodeQL (avgustinov2016ql, ; szabo2023incrementalizing, ). Each participant answered eight questions per task (16 total) designed to reflect a realistic sensemaking process of tool-generated warnings They had to answer questions about the impact of third-party library models on taint analysis results by identifying pass-through APIs calls, APIs that serve as sanitizers, etc. Participants increased the average question completion rate from 71% with the baseline CodeQL visualizer to 92% with WhyFlow, especially on the kinds of questions that require global reasoning of multiple taint flows. When using WhyFlow, users were more accurate in identifying multiple sinks affected by the same third-party library model (50% vs. 17%).
@@
We measured cognitive load via NASA-TLX questions. We found that WhyFlow led to improvement on the participants’ self-reported mental demand, stress, and success rate. Considering the technology acceptance model (lee2003technology, ), participants rated WhyFlow higher on both Confidence (mean 4.3 vs. 2.1) and Ease of Use (4.3 vs. 1.9) than the baseline CodeQL. Qualitative analysis highlighted that color-coded graphs and template queries helped alleviate analysis paralysis (schwartz2015paradox, ). Overall, these results underscore that WhyFlow substantially enhances users’ ability to make sense of complex taint paths and reduce their mental workload through QA-based debugging.
@@
    Configuring taint analysis is tricky, often leading to unexpected results. To help users understand and fix these issues, WhyFlow is the first interactive end-user debugger that shows how different user configuration choices and modeling of third-party libraries impact the taint analysis results.
@@
    WhyFlow is equipped with template queries for ‘why,’ ‘why not,’ and ‘what if’ questions, which are automatically translated into logic queries; it enables users to concretize the template queries with concrete code names embedded in tool-generated warnings. WhyFlow makes it easier to examine the global impact of models, which are difficult to reason when viewing warnings in a list view, packaged with an existing taint analyzer.
@@
    Improvement in accuracy, reduction in cognitive load, improvement in confidence, and ease of use compared to CodeQL’s visualizer is statistically significant. The additional follow-up study with two professionals corroborates this improvement in accuracy, confidence, and ease of use.
The remainder of this paper is organized as follows. Section 2 introduces a motivating example and the design goals. Section 3 presents our tool, WhyFlow. Section 4 provides the study design. Section 5 reports the user study’s results. Section 6 discusses the implications of our findings and threats to validity. Section 8 presents related work. Section 9 concludes our paper.
@@ 2. Motivation
When using modern static taint analyzers such as CodeQL (szabo2023incrementalizing, ), developers struggle to identify why unexpected flows appear or why certain expected flows are missing. These tools rely on models of external libraries, which may not always match the developer’s expectations. This paper is concerned with this problem of end-user debugging of taint analysis—i.e., understanding the impact of user configuration choices. Two key end-user debugging challenges prevent developers from easily making sense of unexpected results.
First, traditional tools lack support for inquiry-based debugging, presenting only isolated warnings without contextual flow connections. This forces developers into manual tracing, which hinders their ability to answer questions they have during debugging.
Second, beyond isolated flow views, developers need support for reasoning about the model’s impact on multiple dataflows. Current tools lack the ability to predict the ripple effects of model changes, forcing developers to manually piece together disparate flows.
Prior work introduced WhyLine (ko2004designing, ) to let users ask “why did” or “why didn’t” questions on a program trace, for debugging runtime errors; however, “why and why-not” questions remain unaddressed in the context of modern taint analysis. Developers want customizability over taint analysis (nachtigall2022large, ), visual outputs of warnings (johnson2013don, ), and the ability to inspect intermediate pass-through nodes of the analysis (ko2004designing, ). Yet, typical dataflow or taint analysis interfaces do not enable higher-level inquiries such as “Why does data from source X reach sink Y?” or “Why is no warning raised for a known bug?
@@
AffectedSinks “If we alter a third-party library $Z$’s model, which sinks are affected?” “Under a new assumption of treating a third-party library $Z$ as a sanitizer, which previously reported sinks are no longer reachable from XX?”
DivergentSinks “Which third-party library $X$’s model could influence multiple taint flows from the same source XX?” “What are the common third-party API nodes in multiple paths originating from $X$ that split into multiple different sinks?”
DivergentSources “Which third-party library $X$’s model could influence multiple taint flows reaching the same sink YY?” “What are the common third-party API nodes in multiple flow paths that eventually reach the same sink YY?”
GlobalImpact “Which third-party library $Z$’s model could have the largest global influence on dataflows from $X$ to YY?” “What is the frequency of each third-party API call appearing along all paths from $X$ to $Y$ in terms of the overall usage counts?”
@@
1. WhyFlow: Why is there a taint flow from a source X to a sink Y? Suppose that Alice sees an unexpected taint-analysis warning in CodeQL from a source user.getSSN() (untrusted data) to a sink log(SSN) (potential vulnerability) (ko2004designing, ; nachtigall2022large, ; johnson2013don, ). She suspects the culprit is an imprecise third-party library model; however, manually tracing a long chain of calls is overwhelming. WhyFlow query highlights the taint flow path from a source in green to a sink in red, displaying pass-through third-party API calls in orange, shown in Figure 2. Alice notices an intermediate API call encrypt(SSN). She then discovers that this API is a sanitizer, and thus this warning should not have been reported and encrypt()’s model should be debugged.
@@
2. WhyNotFlow: Why is there no taint flow from a source X to a sink Y? Suppose that Alice needs to investigate a missing taint flow (i.e., a bug arises at a sink, yet CodeQL does not issue a corresponding warning). She suspects that a third-party library is mistakenly modeled as a sanitizer. In Figure 3, the sensitive data travels from user.getSSN() through an external API: format(SSN) to log(SSN). After investigating the intermediate flow steps, Alice finds that format(SSN) was erroneously modeled as a sanitizer.
WhyNotFlow visually identifies which APIs are acting as sanitizers (dashed arrows), pinpointing which API model could be responsible for killing a flow. WhyNotFlow performs a speculative analysis by reasoning which taint flow path could have been plausible under a configuration where a sanitizer is instead modeled as a non-sanitizer. In Figure 3, the graphical view marks the arrow after the node for format(SSN) with a dashed line, indicating the flow currently does not exist, but would exist with a different model assumption.
@@
3. AffectedSinks: If we alter a third-party library $Z$’s model, which sinks are affected? Alice wants to reason about the global impact of updating a third-party library model (smith2015questions, ). Suppose that she sees an unexpected warning and considers marking a third-party library’s API as a sanitizer. However, she is concerned this update might suppress other warnings. Without WhyFlow, she would need to sift through all reported warnings, manually trace each flow from the source, and check which sinks might become unreachable after her change—a time-consuming process. AffectedSinks automates this what-if analysis, immediately revealing all red sinks that would be “killed” (Figure 4).
@@
Figure 4. AffectedSinks: “If we alter a third-party library’s model, which sinks are affected?
@@
4. DivergentSinks & 5. DivergentSources: Which third-party library model could influence multiple taint flows reaching the same sink (or originating from the same source)?
@@
Suppose that Alice would like to know whether a known vulnerability reaching multiple sinks could be fixed at once (smith2015questions, ; piskachev2019codebase, ). In a typical taint analyzer, it can be cumbersome to trace how a single piece of sensitive data, such as an SSN, propagates through multiple taint flows. Consequently, she may prefer to identify a common interception point instead of fixing one sink at a time. With DivergentSinks query, she can quickly locate a common point from the source that “splits” into multiple sinks. (see Figure 5).
@@
Figure 5. DivergentSinks: “Which third-party library model could influence multiple taint flows from the same source?”
@@
To illustrate another scenario, imagine a single sensitive destination (sink) that receives data from multiple untrusted sources. Using the DivergentSources query, Alice can readily identify if these diverse sources converge towards the same vulnerable point. This allows her to pinpoint the specific model responsible for the convergence, i.e., the ‘culprit model.’ Unlike traditional list-based warning views, WhyFlow visually highlights where taint paths intersect, making it much easier to pinpoint the culprit model.
@@
6. GlobalImpact: Which third-party library’s model could have the largest global influence on multiple flows from a source to a sink? When multiple models could explain a false or missing warning, developers like Alice often prefer a conservative fix that impacts the fewest flows (murphy2013design, ). Without WhyFlow, identifying the frequency of each API in all taint paths would require Alice to painstakingly review every single warning and count occurrences manually. GlobalImpact automatically computes how often each API appears across multiple paths, ranking them by frequency. In Figure 6, because extractParts(standardizedSSN) appears in more paths, it has a higher score, signaling a larger global impact, which provides Alice with a better way to view the global impact of the APIs existing in the flows she is examining.
@@
Figure 6. GlobalImpact: “Which third-party library model could have the largest global influence on dataflows from a source to a sink?”